OpenStudy (anonymous):

The number on two consecutively numbered gym lockers have a sum of159. what are the locker numbers. Use comma to separate the answer

OpenStudy (lgbasallote):

let x = first locker x+1 = second locker x + x + 1 = 159 2x + 1 = 159 2x = 158 x = 79 x+1 = 80

Directrix (directrix):

Let the locker numbers be x and x+1. Then, x + (x+1) = 159 2x = 158 x = 79 x=1 = 80 (79 , 80) ---> ordered pair

OpenStudy (anonymous):

About 50.4 billion pieces of unopened junk mail ends up in landfills each year. This is about 48% of all junk mail that is sent annually. How many pieces of junk mail are sent annually?